Suppose u and v are nonzero vectors in 3-space, where u=(u1,u2, u3) and v=(v1,v2,v3). Prove that u × v is perpendicular to both u and v by making use of the dot product
1
Expert's answer
2020-05-08T19:45:07-0400

We have "u\times v=\begin{vmatrix}
u_1&u_2&u_3\\
v_1&v_2&v_3\\
e_1&e_2&e_3
\end{vmatrix}="

"=\left(\begin{vmatrix}
u_2&u_3\\
v_2&v_3\\
\end{vmatrix},-\begin{vmatrix}
u_1&u_3\\
v_1&v_3\\
\end{vmatrix},\begin{vmatrix}
u_1&u_2\\
v_1&v_2\\
\end{vmatrix}\right)"

Then "(u,u\times v)=u_1\begin{vmatrix}
u_2&u_3\\
v_2&v_3\\
\end{vmatrix}-u_2\begin{vmatrix}
u_1&u_3\\
v_1&v_3\\
\end{vmatrix}+u_3\begin{vmatrix}
u_1&u_2\\
v_1&v_2\\
\end{vmatrix}="

"=\begin{vmatrix}
u_1&u_2&u_3\\
v_1&v_2&v_3\\
u_1&u_2&u_3
\end{vmatrix}=0" and "(v,u\times v)=v_1\begin{vmatrix}
u_2&u_3\\
v_2&v_3\\
\end{vmatrix}-v_2\begin{vmatrix}
u_1&u_3\\
v_1&v_3\\
\end{vmatrix}+v_3\begin{vmatrix}
u_1&u_2\\
v_1&v_2\\
\end{vmatrix}="

"=\begin{vmatrix}
u_1&u_2&u_3\\
v_1&v_2&v_3\\
v_1&v_2&v_3
\end{vmatrix}=0"